Standard Product Cost, Direct Materials Variance Condiments Company uses standards to control its materials costs. Assume that a batch of ketchup (2,600 pounds) has the following standards: Standard Quantity Standard Price Whole tomatoes 4,300 lbs. $ 0.37 per lb. Vinegar 240 gal. $ 2.30 per gal. Corn syrup 21 gal. $ 8.20 per gal. Salt 96 lbs. $ 2.10 per lb. The actual materials in a batch may vary from the standard due to tomato characteristics. Assume that the actual quantities of materials for batch K-111 were as follows: 4,500 lbs. of tomatoes 230 gal. of vinegar 22 gal. of corn syrup 95 lbs. of salt a. Determine the standard unit materials cost per pound for a standard batch. If required, round amounts to the nearest cent. Ingredient Standard Cost per Batch Whole tomatoes $ Vinegar $ Corn syrup $ Salt $ Total $ Standard unit materials cost per pound $ b. Determine the direct materials quantity variance for batch K-111. If required, round amounts to the nearest cent. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Ingredient Materials Quantity Variance Favorable/Unfavorable Whole tomatoes $ Vinegar $ Corn syrup $ Salt $ Total direct materials quantity variance $
Answer #1

1) Calculate standard Unit material cost per pound

Standard cost
Whole tomatoes (4300*.37) 1591
Vinegar (240*2.3) 552
Corn syrup (21*8.2) 172.2
Salt (96*2.1) 201.60
Total Standard cost 2516.80
Quantity 2600
Standard unit material cost per pound 0.968 or 0.97

2) Calculate direct material quantity variance

Direct material quantity variance
Whole tomatoes (4300-4500)*0.37 = 74 U
Vinegar (240-230)*2.3 = -23 F
Corn syrup (21-22)*8.2 = 8.2 U
Salt (96-95)*2.1 = -2.1 F
Total Material quantity variance 57.10 U